If bc:ac:ab=1:3:5, then find a/bc : b/ca.
Answer by satyareddy22(84)

bc:ac:ab=1:3:5
bc/ac=1/3=>b/a=1/3=>a/b=3/1=(3*5)/(1*5)=15/5
ac/ab=3/5=>c/b=3/5=>b/c=5/3
=>a:b:c=15:5:3
a/bc=15/1
b/ca=5/3
a/bc : b/ca=(15/1)/(5/3)=(15/1)*(3/5)=3*3=9
Therefore a/bc : b/ca=9:1